# Assessing subhalo finders in cosmological hydrodynamical simulations

### v1

@@ 1 Introduction
In the past, this approach helped rule out hot models of dark matter (White et al., 1983), as well as hint at a non-zero cosmological constant (e.g. Efstathiou et al., 1990). Ongoing surveys aim to constrain key cosmological parameters, such as the amplitude of power fluctuations (e.g. via weak lensing; Euclid Collaboration et al. 2024a), or the nature of dark matter (e.g. using Milky Way satellites; Drlica-Wagner et al. 2019). Making such cosmological inferences requires forward modelling the evolution of structure in the Universe under different sets of parameter choices. This in turn involves running cosmological simulations of increasingly large volumes and higher resolutions. However, simulations follow individual mass elements, rather than the conglomerate structures that they are a part of. As such, an important step in this process is finding where the haloes and the galaxies that they harbour are located.
@@ 2 Simulations
The identifiers of the three resolution levels of the hydrodynamical simulations we analyse, and their corresponding dark matter particle mass ($m_{\mathrm{dm}}$), baryonic particle mass ($m_{\mathrm{b}}$) and the maximum physical softening length ($\epsilon$) are: L1_m10 ($m_{\mathrm{dm}}=4.52\times 10^{10}\,\mathrm{M}_{\odot}$, $m_{\mathrm{b}}=8.56\times 10^{9}\,\mathrm{M}_{\odot}$, $\epsilon_{\mathrm{p}}=11.40\,\mathrm{kpc}$), L1_m9 ($m_{\mathrm{dm}}=5.65\times 10^{9}\,\mathrm{M}_{\odot}$, $m_{\mathrm{b}}=1.07\times 10^{9}\,\mathrm{M}_{\odot}$, $\epsilon_{\mathrm{p}}=5.70\,\mathrm{kpc}$) and L1_m8 ($m_{\mathrm{dm}}=7.06\times 10^{8}\,\mathrm{M}_{\odot}$, $m_{\mathrm{b}}=1.34\times 10^{8}\,\mathrm{M}_{\odot}$, $\epsilon_{\mathrm{p}}=2.85\,\mathrm{kpc}$). The particles in the corresponding DMO versions of the simulations have masses: L1_m10_DMO ($m_{\mathrm{dm}}=5.38\times 10^{10}\,\mathrm{M}_{\odot}$), L1_m9_DMO ($m_{\mathrm{dm}}=6.72\times 10^{9}\,\mathrm{M}_{\odot}$) and L1_m8_DMO ($m_{\mathrm{dm}}=8.40\times 10^{8}\,\mathrm{M}_{\odot}$). From hereon, we omit the _DMO suffix, as we plot the results of hydrodynamical and DMO simulations in separate panels when applicable. The simulations have 79 outputs, with an average spacing in expansion factor of $\langle\Delta\ln a\rangle=0.015$. The output strategy was not tuned to work with the HBT-HERONS subhalo finder.
@@ 3.2 HBT-HERONS
The HBT-HERONS [^1] Available at https://github.com/SWIFTSIM/HBT-HERONS (Hierarchical Bound Tracing - Hydro-Enabled Retrieval of Objects in Numerical Simulations) algorithm is an updated version of the history-based subhalo finder HBT+ (Han et al., 2018). This new version collects several new additions that improve the identification and tracking of subhaloes in both DMO and hydrodynamical simulations, which are presented in detail in Appendix A [^2] The HBT-HERONS catalogues for FLAMINGO were made without including the centering change described in §A.7. The need for that modification only became clear when running on higher resolution simulations.. We additionally modified the way in which particles are partitioned within the code to speed it up. This allowed us to run HBT-HERONS using 256 nodes (128 cores per node) on the FLAMINGO-10k DMO simulation (Pizzati et al. 2024; Schaller et al. in prep), which contains over $10080^{3}$ ($\approx 10^{12}$) particles. We also used it to analyse the largest hydrodynamical simulation run to $z=0$ to date, the 2.8 Gpc FLAMINGO box with $2\times 5040^{3}$ particles. Given its small computational cost and robust identification of structures, both at fixed and across time (Chandro-Gómez et al., 2025), we make HBT-HERONS our fiducial choice of subhalo finder. In this subsection, we present a general overview of how HBT-HERONS works, with an accompanying diagram to illustrate important steps in Fig. 1.
@@
HBT-HERONS assumes that structure forms hierarchically, i.e. that every present-day subhalo has been a central subhalo in the past. In practice, this means that particles belonging to a given halo become associated to its central subhalo. These memberships are subsequently used to identify the subhalo if it is accreted as a satellite later on. This approach circumvents the difficulties involved when trying to separate heavily overlapping subhaloes. Variations of this history-based approach have recently been used to augment other structure finding algorithms, e.g. Subfind-HBT [^3] Despite the similarity in the name, the implementation details differ from those in HBT+. This leads to differences that we discuss in Appendix B. (Springel et al., 2021), MORIA (Diemer et al., 2024) and SYMFIND (Mansfield et al., 2024).
@@
The algorithm requires a series of simulation outputs saved at sufficiently fine time spacing [^4] Han et al. (2012) found that using a time spacing of $\Delta\ln a=0.2$ leads to a bound subhalo mass function whose amplitude is 5% of the $z=2.1$ value obtained using a finer time spacing. The convergence improves at lower redshifts and with more frequent outputs. We independently confirmed that FLAMINGO has sufficient outputs to provide ”time”-converged mass functions, as there are 79 available outputs with an average spacing of $\langle\Delta\ln a\rangle=0.015$., each containing information on the phase-space distribution and FoF group memberships of particles. Each output is then analysed consecutively, from early to late times. As structures form, they are assigned unique IDs used throughout the remainder of the simulation. The time-persistence of their IDs facilitates the retrieval of the evolution of the subhalo along its main progenitor branch. The steps each output is subject to are explained below.
@@ 3.2.1 Assigning host haloes
First, FoF groups without any assigned subhaloes, and hence no previously existing structures associated to it, are considered potential seeds for new subhaloes. In other words, all of the particles belonging to that FoF group comprise a candidate subhalo. If the candidate subhalo is identified as self-bound, a new central subhalo with a unique ID (TrackID) is created.
@@ 3.2.2 Identification of the central subhalo
During the assignment of host FoF groups, two or more subhaloes that were centrals in the previous output can be found within the same host halo. This indicates an accretion event has occurred, and a choice about which subhalo is the central needs to be made. Making a sensible choice is important, as it determines which subhalo acquires the particles that have been accreted by its host FoF and that do not already belong to satellites, i.e. diffuse mass accretion (at a given resolution).
@@
If more than one subhalo satisfies the above mass threshold, the specific orbital kinetic energy of each candidate is computed in the centre of mass reference frame of the host FoF group. The subhalo candidate with the lowest value, and hence moving most like its host FoF group, is chosen as central. Updating which subhaloes are assigned to be satellites also propagates to the hierarchy of structures that they brought with them, e.g. their satellites become satellites of satellites.
3.2.3 Mass accretion
The evolution of central and satellite subhaloes is manifestly different. Centrals generally grow in mass, whereas the growth of satellites typically halts or becomes negative after being accreted. HBT-HERONS therefore only allows the centrals of each FoF group to accrete new mass in diffuse form. Satellites can still accrete mass, but only if the mass has been stripped from one of their own satellites. Conversely, although for centrals the acquisition of particles is generally limited to those which are not already associated to resolved satellites, mass can still be accreted from satellites if the particles are not bound to them, as explained in §3.2.4.
An important aspect of the tracking of which particles belong to which subhalo is the notion of the ‘source’ subhalo (see Han et al. 2012). This is the initial set of particles passed to the unbinding routine when checking for its self-boundness. The way in which it is defined differs across subhalo finders, and is closely related to the performance with which they identify subhaloes. Taking Subfind as an example, the source subhalo corresponds to all particles enclosed within a density peak in configuration space, excluding additional density peaks within. HBT-HERONS defines it based on the past evolution of subhaloes.
For the simplest case of a single (central) subhalo in a FoF group, all of the particles in the FoF group are initially part of its source, modulo the truncation that occurs after unbinding (see §3.2.4). The source particles belonging to said subhalo are tracked forward in time, so that if it is subsequently found as a satellite, they are used to separate it from the background particles belonging to the more massive subhalo. In this configuration, the central has as its source all the particles that are part of the FoF, except for those that already belong to the sources of resolved satellite subhaloes.
3.2.4 Unbinding
@@
The hierarchy of structures within each FoF group is analysed recursively from the bottom of the subhalo hierarchy. In other words, the subhaloes without any children are subject to unbinding first, at which point their parents can be analysed. The process is repeated until the central subhalo of the FoF is analysed. This depth-first approach allows for the possibility of parents gaining bound mass that is not bound to its children subhaloes.
@@
The unbinding of a single subhalo is done iteratively, whereby its centre of mass velocity and position is updated based on the particles deemed to be bound in any given iteration. To speed up the process of unbinding, the particle distribution is subsampled if the subhalo contains more than $N_{\mathrm{subsample}}$ (1000 by default) particles. In practice, this can lead to small differences in the subhalo bound mass functions at the high mass end between different runs of HBT-HERONS on the same simulation. This reflects different random numbers being generated when doing the particle subsampling. Additionally, the threshold used to determine whether the iterative unbinding has converged can be changed to reduce the number of iterations. The threshold used to stop the iterations is defined based on the ratio of bound particles ($N_{\mathrm{bound}}$) between two consecutive iterations: $N^{i+1}_{\mathrm{bound}}/N^{i}_{\mathrm{bound}}\geq f_{\mathrm{converge}}$, where $f_{\mathrm{converge}}=0.995$ by default.
@@
As mentioned previously, each satellite subhalo has its own source subhalo. Despite the possibility of accreting mass from its own children subhaloes, more bound mass is typically lost than gained. Therefore, the source component is generally larger than the set of bound particles. Although tracking the particles in this manner is crucial to separate satellite subhaloes from the background of their host, without adequate care it can lead to failures in subhalo finding. For example, if the source subhalo is dominated by particles that are truly unbound, the centre of mass estimate can be offset from the true centre of the subhalo. Consequently, the kinetic energies in its core will be biased high, as the Hubble flow contribution (proportional to the distance to the centre of mass) is larger. This can lead to an apparently unbound core, even though it is purely driven by an incorrect choice of reference frame.
@@
To prevent such instabilities during unbinding, HBT-HERONS limits the number of particles that any source subhalo can have ($N^{\mathrm{max}}_{\mathrm{source}}$). By default, it equals three times the number of bound particles, i.e. $N^{\mathrm{max}}_{\mathrm{source}}=3N_{\mathrm{bound}}$. After the unbinding step, the least bound particles of the source subhalo are removed until $N_{\mathrm{source}}\leq N^{\mathrm{max}}_{\mathrm{source}}$. If possible, the particles removed from the source subhalo are accreted by the source of its parent subhalo.
@@
Lastly, if a previously resolved subhalo no longer has enough bound particles to be above the threshold to be classified as self-bound ($N^{\mathrm{min}}_{\mathrm{bound}}$; 20 by default) or has insufficient tracer particle types ($N^{\mathrm{min}}_{\mathrm{tracer}}$; 10 by default), it becomes an orphan. The position and velocity of orphan subhaloes, which are often used in the context of semi-analytical models of galaxy formation (e.g. Kitzbichler & White, 2008; Guo & White, 2014), are subsequently tracked by their most bound tracer particle when the subhalo was last resolved. Note that if the tracer particle is found in a FoF host without currently resolved subhaloes, the particles are acquired by the source of the orphan subhalo. This means that if they are found to be bound, orphans can re-appear in the simulation.

### v3

@@ 1 Introduction
In the past, this approach helped rule out hot models of dark matter (White et al., 1983), as well as hint at a non-zero cosmological constant (e.g. Efstathiou et al., 1990). Ongoing surveys aim to constrain key cosmological parameters, such as the amplitude of power fluctuations (e.g. via weak lensing; Euclid Collaboration et al. 2025a), or the nature of dark matter (e.g. using Milky Way satellites; Drlica-Wagner et al. 2019). Making such cosmological inferences requires forward modelling the evolution of structure in the Universe under different sets of parameter choices. This in turn involves running cosmological simulations of increasingly large volumes and higher resolutions. However, simulations follow individual mass elements, rather than the conglomerate structures that they are a part of. As such, an important step in this process is finding where the haloes and the galaxies that they harbour are located.
@@ 2 Simulations
The identifiers of the three resolution levels of the cosmological simulations we analyse, and their corresponding dark matter particle mass ($m_{\mathrm{dm}}$), baryonic particle mass ($m_{\mathrm{b}}$) and the maximum physical softening length ($\epsilon$) are listed in Table 1. The simulations have 79 outputs, with an average spacing in expansion factor of $\langle\Delta\ln a\rangle=0.015$. The output strategy was not tuned to work with the HBT-HERONS subhalo finder.
Table 1: Summary of the initial mass for dark matter ($m_{\mathrm{dm}}$) and baryonic ($m_{\mathrm{b}}$) particles in the cosmological simulation suite we use in this work, as well as the maximum physical softening length ($\epsilon$). Note that we do not use the hydrodynamical version of the L1_m8 simulation due to the large computational cost required to analyse it with all of the subhalo finders we compare.
Simulation Identifier Type $m_{\mathrm{DM}}$ $\mathrm{M}_{\odot}$] $m_{\mathrm{bar}}$ $\mathrm{M}_{\odot}$] $\epsilon$ [kpc]
L1_m10 DMO $5.38\times 10^{10}$ - 11.40
Hydro $4.52\times 10^{10}$ $8.56\times 10^{9}$ 11.40
L1_m9 DMO $6.72\times 10^{9}$ - 5.70
Hydro $5.65\times 10^{9}$ $1.07\times 10^{9}$ 5.70
L1_m8 DMO $8.40\times 10^{8}$ - 2.85
Hydro $7.06\times 10^{8}$ $1.34\times 10^{8}$ 2.85
@@ 3.2 HBT-HERONS
The HBT-HERONS [^1] Available at https://github.com/SWIFTSIM/HBT-HERONS (Hierarchical Bound Tracing - Hydro-Enabled Retrieval of Objects in Numerical Simulations) algorithm is an updated version of the history-based subhalo finder HBT+ (Han et al., 2018). This new version collects several new additions that improve the identification and tracking of subhaloes in both DMO and hydrodynamical simulations, which are presented in detail in Appendix A [^2] The HBT-HERONS catalogues for FLAMINGO were made without including the symmetric merging criterion (§A.3), the center refinement change described in §A.7 and the gas re-attachtment step (§A.8). The need for the last two modifications only became clear when running on higher resolution simulations. We have verified that the effect of these changes is negligible for all of the results presented here.. We additionally modified the way in which particles are partitioned within the code to speed it up. This allowed us to run HBT-HERONS using 256 nodes (128 cores per node) on the FLAMINGO-10k DMO simulation (Pizzati et al. 2024; Schaller et al. in prep), which contains over $10080^{3}$ ($\approx 10^{12}$) particles. We also used it to analyse the largest hydrodynamical simulation run to $z=0$ to date, the 2.8 Gpc FLAMINGO box with $2\times 5040^{3}$ particles. Given its small computational cost and robust identification of structures, both at fixed and across time (Chandro-Gómez et al., 2025), we make HBT-HERONS our fiducial choice of subhalo finder. In this subsection, we present a general overview of how HBT-HERONS works, with an accompanying diagram to illustrate important steps in Fig. 1.
HBT-HERONS assumes that structure forms hierarchically, i.e. that every present-day subhalo has been a central subhalo in the past. In practice, this means that particles belonging to a given halo become associated to its central subhalo in what is called a ‘source’ subhalo (see Han et al. 2012). The source subhalo is the set of particles that constitute a tentative subhalo within the simulation, whose existence needs confirming after a ‘cleaning’ step to identify whether it is indeed a physical subhalo or just spurious noise. Our choice to determine when a subhalo is resolved is the same as the choice made by Subfind and VELOCIraptor, which is to check whether the subhalo contains a minimum number of self-bound particles (§3.2.4). However, alternative criteria used to deem a subhalo as physical exist, like requiring a minimum fraction of self-bound particles in a given 6D FoF subhalo in ROCKSTAR or by discarding subhaloes whose concentrations are non-physical in Bloodhound (Kong et al., 2025).
@@
The way in which the source subhalo is defined differs across subhalo finders, and it is closely related to the performance with which they are able to identify subhaloes. For example, the source subhalo in Subfind corresponds to all particles enclosed within a density peak in configuration space, excluding additional density peaks within. In contrast, HBT-HERONS defines it based on the past evolution of subhaloes: the source subhalo of satellite subhaloes is built from the source subhalo when they were last a central, excluding its resolved substructure. This approach circumvents the difficulties involved when trying to separate heavily overlapping subhaloes in phase-space. Indeed, variations of this history-based approach have recently been used to augment other structure finding algorithms, e.g. Subfind-HBT [^3] Despite the similarity in the name, the implementation details differ from those in HBT+. This leads to differences that we discuss in Appendix B. (Springel et al., 2021), ROCKSTAR+MORIA (Diemer et al., 2024), ROCKSTAR+SYMFIND (Mansfield et al., 2024) and ROCKSTAR+Bloodhound (Kong et al., 2025).
@@
HBT-HERONS requires a series of simulation outputs saved at sufficiently fine time spacing [^4] Han et al. (2012) found that using a time spacing of $\Delta\ln a=0.2$ leads to a bound subhalo mass function whose amplitude is 5% of the $z=2.1$ value obtained using a finer time spacing. The convergence improves at lower redshifts and with more frequent outputs. We independently confirmed that FLAMINGO has sufficient outputs to provide ”time”-converged mass functions, as there are 79 available outputs with an average spacing of $\langle\Delta\ln a\rangle=0.015$., each containing information on the phase-space distribution and FoF group memberships of particles. Each output is then analysed consecutively, from early to late times. As structures form, they are assigned unique IDs used throughout the remainder of the simulation, whose time-persistence facilitates the retrieval of the evolution of the subhalo along its main progenitor branch. At each output, every pre-existing subhalo is assigned to a given FoF group (§3.2.1), and a central subhalo is chosen per FoF group based on a mass and orbital kinetic energy criterion (§3.2.2). The classification of whether a subhalo is a central or not determines which physical mechanisms are eligible to affect its mass from one output to the next (§3.2.3, §3.2.4, §3.2.5 and §3.2.6), which reflects the expected evolutionary differences between central and satellite subhaloes.
@@ 3.2.1 Assigning host haloes
First, FoF groups without any assigned subhaloes, and hence no previously existing structures associated to it, are considered potential seeds for new subhaloes. In other words, all of the particles belonging to that FoF group are grouped into a single source subhalo. If the candidate subhalo is identified as self-bound, a new central subhalo with a unique ID (TrackID) is created.
@@ 3.2.2 Identification of the central subhalo
During the assignment of host FoF groups, two or more subhaloes that were centrals in the previous output can be found within the same host halo. This indicates an accretion event has occurred, and a choice about which subhalo is the central needs to be made. Making a sensible choice is important, as it determines which subhalo acquires the particles that have been accreted by its host FoF and that do not already belong to satellites, i.e. diffuse mass accretion at a given resolution (see §3.2.3).
@@
If more than one subhalo satisfies the above mass threshold, the specific orbital kinetic energy of each candidate is computed in the centre of mass reference frame of the host FoF group. The subhalo candidate with the lowest value, and hence moving most like its host FoF group, is chosen as central [^5] Note that if the subhalo chosen to be the central is not the most massive subhalo in the FoF group after unbinding (§3.2.4), the most massive subhalo of the group is changed to be the central. However, diffuse mass accretion (§3.2.3) at that output will reflect the original central choice.. Updating which subhaloes are assigned to be satellites also propagates to the hierarchy of structures that they brought with them, e.g. their satellites become satellites of satellites.
We note that sensible values for $f_{\rm major}$ are of $\mathcal{O}(0.1)$. Values much lower than 0.1 make low mass subhaloes become candidate centrals. As they dominate in number over massive subhaloes, and the central is chosen out of the candidates based on their specific orbital kinetic energy, it is very likely that a low-mass subhalo is incorrectly chosen as a central instead of a more suitable massive subhalo. Consequently, large mass fluctuations may occur, as a low-mass subhalo suddenly acquires most of the mass in the FoF group. On the other hand, using $f_{\rm major}=1$ is too restrictive, as only the most massive subhalo in the previous output is a candidate central. Since mass is assigned in an exclusive manner, it then becomes a matter of which subhaloes have a more massive satellite population prior to being in the same FoF, even if they have similar spherical overdensity masses.
3.2.3 Diffuse mass accretion
We make the physically-motivated assumption that only central subhaloes can grow in mass through the accretion of surrounding diffuse mass, as the evolution of satellites is more constrained and most will have ceased growing or are even losing mass. Consequently, the source subhalo of central subhaloes can grow by accreting particles that have been newly acquired by their host FoF group, and which are not already associated to the source subhalo of resolved subhaloes. For the simplest case of a single (central) subhalo in a FoF group, all of the particles in the FoF group are initially part of its source.
3.2.4 Unbinding and mass stripping
@@
The hierarchy of structures within each FoF group is analysed recursively from the bottom of the subhalo hierarchy. In other words, the subhaloes without any children are subject to unbinding first, at which point their parents can be analysed. The process is repeated until the central subhalo of the FoF is analysed. This depth-first [^6] In HBT-HERONS, depth refers to the number of hierarchical connections that a given subhalo is away from the central subhalo of its host FoF group. A central subhalo thus has a depth of 0, its satellites a depth of 1, the satellites of satellites a depth of 2, etc. approach allows for the possibility of parents gaining bound mass that is not bound to its children subhaloes. In other words, any central or satellite subhalo can accrete mass that is stripped from the subhaloes in their own satellite system.
The unbinding of a single subhalo is done iteratively, whereby its centre of mass position and velocity are updated based on the particles deemed to be bound in any given iteration. To speed up the process of unbinding, which is generally the most expensive step in the subhalo finding process, the particle distribution is subsampled if the subhalo contains more than $N_{\mathrm{subsample}}$ particles. By default, $N_{\mathrm{subsample}}$ = 1000, which provides results very close to not subsampling the particles at all whilst reducing the computational cost (by a factor of $\approx 10$ when using the default parameters for the L1_m9 simulation). While the exact value of $N_{\mathrm{subsample}}$ is user-adjustable, we caution against using values lower than the default, as the code becomes increasingly likely to choose a subset of particles that do not represent the true mass distribution of the subhalo, leading to its premature disruption. Once the random subsample of particles is chosen, the particle masses are appropriately upscaled to ensure mass conservation, either on a total subhalo mass or a total particle-type mass level. The subsampling of specific particles types can also be disabled, e.g. for black hole particles given their low numbers and already large masses compared to the rest of the particles.
@@
In practice, particle subsampling can lead to a few-percent differences in the subhalo bound mass functions at the high mass end between different runs of HBT-HERONS on the same simulation. This reflects different random numbers being generated when doing the particle subsampling. Additionally, the threshold used to determine whether the iterative unbinding has converged can be changed to reduce the number of iterations. The threshold used to stop the iterations is defined based on the ratio of bound particles ($N_{\mathrm{bound}}$) between two consecutive iterations: $N^{i+1}_{\mathrm{bound}}/N^{i}_{\mathrm{bound}}\geq f_{\mathrm{converge}}$, where $f_{\mathrm{converge}}=0.995$ by default. We recommend using values in the range of $0.9\leq f_{\mathrm{converge}}<1$, as the condition is strict enough to provide similar results to $f_{\mathrm{converge}}=1$, but cuts down on the number of unbinding iterations and therefore computational cost. Using a value that is too small results in too few iterations to have converged bound masses, meaning that some physically unbound particles are misclassified as bound. Since mass is assigned exclusively and in a depth-first manner, satellite subhaloes would then become too massive, while central subhaloes would artificially miss some of their mass.
@@
As mentioned previously, each satellite subhalo has its own source subhalo that reflects the source subhalo it had when it was last a central subhalo. Despite the possibility of accreting mass from its own children subhaloes, more bound mass is typically lost than gained. Therefore, the source component is generally larger than the set of bound particles. Although tracking particles in this manner is crucial to separate satellite subhaloes from the background of their host, it can lead to failures in subhalo finding without adequate care. For example, if the source subhalo is dominated by particles that are truly unbound, the centre of mass estimate can be offset from the true centre of the subhalo. Consequently, the kinetic energies in its core will be biased high, as the Hubble flow contribution (proportional to the distance to the centre of mass) is larger. This can lead to an apparently unbound core, even though it is purely driven by an incorrect choice of reference frame.
@@
To prevent such instabilities during unbinding, HBT-HERONS limits the number of particles that any source subhalo can have ($N^{\mathrm{max}}_{\mathrm{source}}\equiv f^{\mathrm{max}}_{\mathrm{source}}N_{\mathrm{bound}}$). By default, it equals three times the number of bound particles, i.e. $f^{\mathrm{max}}_{\mathrm{source}}=3$. After the unbinding step, the most unbound particles of the source subhalo are removed until $N_{\mathrm{source}}\leq N^{\mathrm{max}}_{\mathrm{source}}$. The particles removed from the source subhalo are accreted by the source of its parent subhalo, if the subhalo has one. The default value we choose is sufficiently large to allow for the re-accretion of particles after they have been momentarily found as unbound. Using too small a value, e.g. $f^{\mathrm{max}}_{\mathrm{source}}\approx 1$, results in the secular loss of mass from satellites due to the continuous removal of particles that are found to be unbound from one snapshot to the next (e.g. as shown for Subfind-HBT in Fig. 21).
Lastly, if a previously resolved subhalo no longer has enough bound particles to be above the threshold to be classified as self-bound ($N^{\mathrm{min}}_{\mathrm{bound}}$; 20 by default) or has an insufficient number of tracer particles ($N^{\mathrm{min}}_{\mathrm{tracer}}$; 10 by default), it is deemed disrupted and becomes an orphan. Based on our choice to determine when subhalos disrupt, we recommend to keep the value of $N^{\rm min}_{\rm bound}$ small, in the range $10\leq N^{\rm min}_{\rm bound}\leq 20$. Higher values result in the ‘disruption’ of subhaloes that are truly physical (i.e. they remain self-bound for many snapshots after their supposed ‘disruption’).
@@
The position and velocity of orphan subhaloes, which are often used in the context of semi-analytical models of galaxy formation (e.g. Kitzbichler & White, 2008; Guo & White, 2014), are subsequently tracked by their most bound tracer particle when the subhalo was last resolved. Note that if the tracer particle is found in a FoF host without currently resolved subhaloes, the particles are acquired by the source of the orphan subhalo. This means that if these particles are found to be bound, orphans can re-appear as fully resolved subhaloes in the simulation.